Household cleaners: as a cleaning agent a solution that had is 28% vinegar is often used. How much pure (100%) vinegar and 10% vinegar must be mixed to obtain a 100 oz of 28% solution
Answer by Boreal(15235)

x ml of 100%
100-x ml of 10%
x*1+.1(100-x)=100*.28
x+10-.1x=28
0.9x=18
x=20 ml of pure ANSWER
100-x=80 ml of 10%
20+8=28 ml of pure total.